Solve system of equations and state whether it is consistent and dependent, consistent and independent or inconsistent:
-3x+9y-3z=-12
-3x+y-z=-1
2x-6y+2z=9

Found 2 solutions by richwmiller, Fombitz:
Answer by richwmiller(17219)

-3x+9y-3z=-12
-3x+y-z=-1
2x-6y+2z=9
-3,9,-3,-12
-3,1,-1,-1
2,-6,2,9

Inconsistent system with no solutions.


Answer by Fombitz(32388)

Dependent and inconsistent, the determinant of the coefficient matrix is zero.